It's been a while and have forgotten the rules.
Solve for X
1/100000 + 1/X = 10000

Found 2 solutions by stanbon, chessace:
Answer by stanbon(75887)

Solve for X
1/100000 + 1/X = 10000
---
1/10^5 + 1/X = 10^4
----
Bultiply thru by 10^5 to get:
1 + 10^5/X = 10^9
----
10^5/X = 10^9-1
----
X/10^5 = 1/(10^9-1)
---
X = 10^5/(10^9-1)
===================
Cheers,
Stan H.

Answer by chessace(471)

Multiply by the denominators (or in general their LCM):
X + 100000 = 1000000000 X
999999999 X = 100000
X = 1 / 9999.99999
My cheap calculator gives the same answer as if the first term (10^-5) isn't there, which is reasonable.